OpenStudy (anonymous):

Solve. Answers should be in Scientific Notation. (3.2 x 10^4) x (1.4 x 10^3) (8.4 x 10^8) / (2.2 x 10^2)

OpenStudy (anonymous):

2.0

OpenStudy (anonymous):

for the first

OpenStudy (anonymous):

Do you solve the actual notation before you compute, or is there a way to compute it without having to do all of that?

OpenStudy (anonymous):

the actual notation

OpenStudy (anonymous):

\[(3.2 \times 10^4) \times (1.4 \times 10^3)\] \[=(3.2\times 1.4)\times 10^{4+3}=4.48\times 10^7\]

OpenStudy (anonymous):

similarly, \[(8.4 \times 10^8) \div (2.2 \times 10^2)=8.4\div 2.2\times 10^{8-2}=3.82\times 10^6\] but \[\frac{8.4}{2.2}=\frac{84}{22}=\frac{42}{11}\] and non repeating non terminating decimal